OpenStudy (anonymous):

find the derivitive of f(x)=9x^6 + 5/2x^3/4 - pi x^2 + 6x - 4 DO NOT SIMPLIFY.

zepdrix (zepdrix):

\[\Large\rm f(x)=9x^6 + \frac{5}{2}x^{3/4} - \pi x^2 + 6x - 4\]Is this what the function looks like?

OpenStudy (anonymous):

yes! @zepdrix

zepdrix (zepdrix):

Mmmm looks like we simply apply the power rule to each term, yes? :o Where is the confusion at?

OpenStudy (anonymous):

do i have to multiply \[\pi(2)\]

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

For that particular term? Yes good. The 2 comes down in front, and the x decreases to a 1 power.\[\Large\rm (\pi x^2)'=2\pi x\]

OpenStudy (anonymous):

\[54x ^{5}+(15\div 8)x ^{-1/4}-2\pi x+6\]

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Mmmm yes, good job!

OpenStudy (anonymous):

thank you! @zepdrix